madiurmom:

A motorcycle can travel 60 miles per gallon. Approximately how many gallons of fuel will the motorcycle need to travel 40 km? (5 points) [1 mile = 1.6 km] 0.04 0.08 0.20 0.42

tiaunabuchan:

its not a or c

madiurmom:

i think its D but im not sure

tiaunabuchan:

yessir

XioGonz:

one gallon= 60 miles then 1/2 gallon= 30 miles 1/2 gallon= 48 Km 1/2=0.5 if we bring it down to 0.42 we 25.2 miles 25.2 miles = 40.32. The question is asking for the approximate number 40.32 is the closest That leaves us with the answer 0.42

madiurmom:

thank you guys :D

XioGonz:

Np :)